UT 20150405 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson (all remote from Tucson)
Observer AO: DMiller, JChristou (LBTO, remote from Tucson)
Telescope Operator: GBechetti
LBTO Support: TSargent (remote from Tucson)

Plan for the night

As per agreement, this night will be converted from engineering to LUCI2 AOCom night:

- ACF-tests in the afternoon

- Continue AOCom

Summary

  • Most of the night lost due to clouds

  • MOS error (introduced by JHeidt)

  • While on sky, poor and variable seeing often prohibited closing the loop

  • God news for the May run(s): Doug and Julian can operate the AO-system

Details

All times are UT
Files are luci2.20150405.NNNN.fits

23:30 Afternoon, ACF-tests started

03:40 ACF-tests done, Geno opens, sky mostly covered with Cirrus, only brightest stars can be seen
T = 4deg, Hum = 25%, low wind

04:00 preset HIP52849, short spectra G210/G40 all orders (DIT=3, NDIT=3), about 4mag extinction,
src, #17, slitview #18/19, verification , N30 LS1, gratings aligned+check, OK, #22, #23-61

heavy cirrus coming in

05:00 M67_RGBstar, focuscheck N30/N3.75, failed, AO-arbitrator, retry, failed again,
DDConnector and Telescope Service restarted, preset OK

05:10 clouds too heavy, stopped

05:36 Sky better, M67_RGBstar again, clouds again, Dave modified script (wrong mag of AO-ref star from NOMAD)
clouds again,

Check homogeneity of illumination by CalUnit, lamp2/H2 #68, guide probe out, #69, lamp3/Ks, #70
inhomogeneous illumination still present (about 40%) and still differs between cameras

07:00 Clouds still present (thicker compared to last night), wind increases

07:40 Better, preset HIP52849, collimated...lost GS,

07:57 Preset A45_R8_0037, seeing gets poorer, about 2" now, clouds, star lost

08:23 A45_R8_0037 again, loop not closed due to arbitrator, 10 times preset repeated....

09:02 stopped/restarted telescope service and data dictionary connector, ....sh...clouds....

09:03 loop finally closed...#71-77

09:20 qso_spec_long, MOS error (panic) => waiting for Peter

=> in the meantime Doug & Julian map the Patrol field of the WUnit stages

Clouds coming in again, seeing variable 0.9-3.5"

LUCI-SW GUI's frozen (starting with ALADIN)

11:37 Still waiting for Peter, but can do some imaging.

11:38 Preset to Titan (for Al) at Airmass 1.7, and try to lock on it (nice exercise for DMiller), LUCI2 will take only images

12:05 Preset to HD193793, seeing too bad for AO-correction..., retried, AO failed, #83-84

Two times an offset within a script was interpreted as preset (once a preset was sent and
AO closed manually) => must be optimized

12:45 Closing up

End of AO-Com#2
Topic revision: r8 - 01 May 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback